which of thefollowing equations represent a horizontal line?
A) 3x-3y=6
B) 6x+3y=6
C) -4x=6
D) 4y=6

Answer by actuary(112)

Write each of the equations in the standard slope - intercept form, y = m*x +b
A) -3y = -3x + 6 (Divide each term of the equation by -3) or y = x - 2
B) 3y = - 6x + 6 (Divide each term of the equation by 3) or y = -2x + 2
C) x = -3/2
D) 4y = 6 or y = 3/2
Equation A is a line with slope 1 and y-intercept of -1 (Not a horizontak line)
Equation B is a line with slope -2 and y-intercept = 2
Equation C is a VERTICAL line at x = -3/2
Equation D is a hoizontal line at y = 3/2